OpenStudy (kitkat16):

need help with parabola equation please One of the x-intercepts of the parabola represented by the equation y = 3x^2 + 6x − 10 is approximately (1.08, 0). The other x-intercept of the parabola is approximately ? need to round to nearest hundred

OpenStudy (kitkat16):

3x^2+6x-10=0 I thought y was -10 but Im so confused. can anyone please help me?

OpenStudy (kitkat16):

hi

OpenStudy (welshfella):

this function will not factor so you can use either completing the square or the quadratic root formula to find the roots

OpenStudy (welshfella):

Hi

OpenStudy (kitkat16):

im not sure how to do that can you show me?

OpenStudy (kitkat16):

0=3x^2+6x-10

OpenStudy (welshfella):

or you could also use the identities A + B = -b/a where A and B are the 2 roots and a and b are the coefficients in the general form ax^2 + bx + c = 0 Have you done that in class yet?

OpenStudy (kitkat16):

yes it looks familar

OpenStudy (kitkat16):

3,6,-10

OpenStudy (kitkat16):

is it quadratic formula

OpenStudy (welshfella):

knowing one root to can put A = 1.08 into the identity so 1.08 + B = -6/3 = -2 B = -3.08

OpenStudy (kitkat16):

3.08 would be x

OpenStudy (kitkat16):

and 0

OpenStudy (welshfella):

No the quadratic formula i referred to was for working out the roots x = [-b +/- sqrt(b^2 - 4ac)] / 2a

OpenStudy (welshfella):

the x intercept would be (-3.08,0)

OpenStudy (kitkat16):

thank you

OpenStudy (kitkat16):

what comes after the -b+

OpenStudy (kitkat16):

that doesn't look right to me

OpenStudy (welshfella):

yw in the textbooks the identity for sum of the roots are usually written as the greek letters alpha + beta = -b/a

OpenStudy (welshfella):

\[\alpha +\beta \]

OpenStudy (kitkat16):

oh yes I've seen that

OpenStudy (kitkat16):

thanks again

OpenStudy (welshfella):

yw